# Pair Production in time-dependent Electric field at Finite times

### v5

@@ Abstract
We investigate the finite-time behavior of pair production from the vacuum by a time-dependent Sauter pulsed electric field. By examining the temporal behavior of the single-particle distribution function, we observe oscillatory patterns in the longitudinal momentum spectrum of the particles at finite times. These oscillations arise due to quantum interference effects resulting from the various dynamical processes/channels leading to the creation of the (quasi-)particle of a given momentum. Furthermore, we derive an approximate and simplified analytical expression for the distribution function at finite times, allowing us to explain these oscillations’ origin and behavior. The role of the vacuum polarization function and its counterterm is also discussed in this regard. The transverse momentum spectrum peaked at the nonzero value of the transverse momentum at finite times, which indicates the role of multiphoton transitions in the creation of quasiparticles.
@@ I Introduction
The concept of pair production in an electromagnetic field has its roots in the mid-1920s after the invention of quantum mechanics, with the formulation of the relativistic wave equation for electrons by Paul Dirac in 1928 Dirac (1928). The Dirac Sea model was proposed to explain the enigma of negative energy solutions. F. Sauter’s work in 1931 demonstrated that strong electric fields can lead to pair creation through tunneling with exponential suppression Sauter (1931). This paved the way for quantum field theory, recognizing the vacuum as a polarizable medium influenced by constant fluctuations. In 1935, W. Heisenberg and H. Euler further explored the peculiarities of the Dirac equation, revealing non-linear modifications in Maxwell’s equations due to the interaction of electromagnetic fields with the electron vacuum loopHeisenberg and Euler (1936). J. Schwinger’s groundbreaking work in 1951 precisely calculated the imaginary part of the one-loop effective Lagrangian in the presence of a static electric field Schwinger (1951). As a result of his seminal work, the phenomenon of vacuum pair creation by electric fields has since become widely known as the Schwinger effect, and it is also famously referred to as the Sauter-Schwinger effect in recognition of F. Sauter’s prior work on solving the Dirac equation in the presence of an electric field Sauter (1932).
@@
Schwinger’s pioneering calculation opened up new avenues of research in quantum field theory and has profoundly impacted our understanding of particle physics in the presence of strong fields. This extraordinary property of quantum vacuum producing spontaneous particle-antiparticle pairs has far-reaching implications for understanding the generation of particle-antiparticle pairs in the presence of a strong electric field Greiner et al. (1985); particle creation in the expanding universeParker (1968); black hole evaporation as a result of Hawking radiation Hawking (1975); Parikh and Wilczek (2000); and Unruh radiation, in which particle production is seen by an accelerating observerUnruh (1976); Kim (2016). The study of generating electron-positron pairs through a spatially constant electrical background field was extended to the electric field with various time dependences. In the 1970s, researchers explored the occurrence of pair production from the vacuum in the presence of an oscillating time-dependent electric field Popov (1971); Marinov and Popov (1977). Their investigations revealed different qualitative behaviors for this process under various interaction regimes. The interaction regimes can be distinguished by the value of the dimensionless Keldysh parameter, $\gamma=\frac{m\omega}{|e|E_{0}}$ Keldysh (1965), where $E_{0}$ is the field amplitude, $\omega$ is the field frequency, $e$ is the electron charge, and $m$ is the electron mass. When $\gamma>>1$, the process probability shows a perturbative power-law scaling with field intensity. Instead, for $\gamma<<1$, it exhibits a manifestly non-perturbative exponential dependence on $\frac{1}{E_{0}}$, similar to the case of a constant electric field Nikishov (1969).
@@
Nevertheless, the study of pair production in strong electric fields subject has attracted sustained interest from theoreticians in recent years due to the extraordinary progress in the development of the ultra-intense lasers technique and the strong field QED experimental studies that are planned at upcoming high-intensity laser facilities, such as the European X-Ray Free-Electron Laser Andruszkow et al. (2000), the Extreme-Light Infrastructure Heinzl and Ilderton (2009); Dunne (2009), the Exawatt Center for Extreme Light Studies Khazanov and et al. (2023). The electric field strength is fast getting closer to the critical value $E_{\text{c}}$.
@@
Various methodologies have been developed to investigate pair production in the presence of strong fields. These include the proper-time method Schwinger (1951); DeWitt (1975), the canonical method Mottola (1985); Tomaras et al. (2001, 2000), Green’s function techniques Gavrilov and Gitman (1996), semiclassical tunneling Parikh and Wilczek (2000), the Schrödinger-functional approach Hallin and Liljenberg (1995), functional techniques Fried and Woodard (2002); Avan et al. (2003); Ilderton (2014),the Wigner formalismZhuang and Heinz (1996); Hebenstreit et al. (2011a, b),mean-field treatments Smolyansky et al. (1998), and worldline instanton methods Affleck et al. (1982); Kim and Page (2002). Studies on pair production in intense laser fields Alkofer et al. (2001); Roberts et al. (2002) have employed analytical and numerical approaches to predict pair-production rates. However, these investigations primarily focus on pair formation at asymptotic times, which alone is insufficient for fully capturing the dynamical aspects of pair production and its time-dependent effects. Several authors have derived expressions for the number of pairs created after a time $T$, where $T$ exceeds the duration of the electric field pulse or corresponds to a moment after the field has vanished Gavrilov and Gitman (1996); Di Piazza (2004); Melike et al. (2012); Ilderton (2022); Adorno et al. (2017). These expressions provide a reliable approximation of the actual pair count.
@@
Understanding pair production dynamics from the initial to final states requires a framework that accounts for transient and non-equilibrium effects. In this context, the distribution function of quasiparticles, governed by the quantum kinetic equation (QKE), offers an effective description of the time-dependent evolution of pair production Blaschke et al. (2015); Smolyansky et al. (1998). Defining the particle distribution function in the quasiparticle basis is both physically meaningful and intuitive. This approach is grounded in the straightforward physics of how a time-dependent field distorts the time-invariance symmetry of the Hamiltonian, leading to mixing between momentum states of positive and negative energies. This method is particularly relevant for exploring the complex dynamical behavior of pair production in spatially uniform, time-dependent electric fields. Specifically, it has been demonstrated that the temporal evolution of the quasiparticle distribution function in momentum space provides insights into the asymptotic states of the quantum field and offers a comprehensive description of the process during the interaction between the matter field and the strong field background Blaschke et al. (2017). It was shown that the distribution function evolves through three distinct temporal stages: the quasi-electron-positron plasma (QEPP) stage, the transient stage, and finally the residual electron-positron plasma (REPP) stage Banerjee and Singh (2022, 2019). In examining the pair production process, both the QEPP stage, where quasi-particles are formed, and the REPP stage, where real particles emerge, are crucial Blaschke:2013ip; Shen:2018lbq. While the Schwinger mechanism is primarily known for producing elementary particles in a vacuum under external fields, it can also generate quasiparticle excitations in exotic materials such as graphene Allor et al. (2008); Zubkov (2012); Gavrilov et al. (2012); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015), which are experimentally more accessible, and in superconductors Solinas et al. (2021). Moreover, while ultracold atoms in optical lattices have been investigated as simulators for the Schwinger mechanism Kasper et al. (2016); Halimeh et al. (2025).
@@
In the present work, we consider the production of electron-positron pairs from the vacuum in a time-varying, spatially uniform pulsed electric field given by $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, with amplitude $E_{0}$ and pulse duration $\tau$. Such background field has received extensive attention in the literature Narozhnyi and Nikishov (1970); Dunne and Hall (1998); Balantekin et al. (1991); Gavrilov and Gitman (1996); Nikishov (2004); Levai and Skokov (2010); Klimchitskaya and Mostepanenko (2013), with a focus on the asymptotic behavior of the probability of pair production. To this end, we study the evolution of the single-particle distribution function, $f(\bm{p},t)$, which is rigorously derived by canonical quantization of the Dirac field and subsequent Bogoliubov transformation to a quasi-particle representation Smolyansky et al. (2017a).
@@
The distribution function exhibits three distinct dynamical stagesBanerjee and Singh (2022); Otto et al. (2016); Nousch et al. (2017); Smolyansky et al. (2017a). By analyzing the temporal evolution of $f(\bm{p},t)$, we find that the emergence and characteristics of these dynamical stages, QEPP, transient, and REPP, are qualitatively and quantitatively influenced by both longitudinal and transverse components of momentum. Next, we analyze the time evolution of the longitudinal momentum spectrum (LMS) of created particles. In the tunneling regime ($\gamma<1$), we observe oscillatory behavior in the LMS at time $t>2\tau$, and this oscillation pattern continuously changes from $t>2\tau$ up to $t<6\tau$. This oscillatory behavior at finite times clearly illustrates the quantum interference effects arising from various dynamical processes and channels contributing to the creation of (quasi-)particles with a given momentum. It can also be understood in terms of the vacuum polarization function $u(\bm{p},t)$ and its counterpart, the depolarization function $v(\bm{p},t)$, which encapsulate the underlying particle production dynamics. We emphasize that the oscillations seen in the LMS are not artifacts but rather possess significant physical relevance. In the multi-photon regime, we observe that LMS at finite times near $t=3\tau$ exhibits a multi-modal structure.
@@
Furthermore, we develop an analytical framework valid for finite times $t>\tau$. Specifically, we derive an analytical expression for the single-particle distribution function as a power series in the small parameter $(1-y)$, where $y=\frac{1}{2}(1+\tanh{(\frac{t}{\tau})})$. The intriguing dynamical features of the momentum distribution function at finite times arise from a function appearing in this expansion. The finite-time behavior of the momentum spectra can be understood in terms of functions that reveal three distinct behaviors, which are discussed in detail in the results section.
@@
Additionally, we explored how the transverse momentum affects LMS and discovered that the oscillatory behavior depends on the value of transverse momentum at that time. Subsequently, we analyzed the temporal progression of the transverse momentum spectrum(TMS) of generated pairs, providing valuable insights into the pair production process. At finite times, the transverse momentum spectra reach their peak at a nonzero value, reflecting the influence of multiphoton transitions in quasiparticle creation.
@@
This article is organized as follows: In Sec.II detailed the theoretical formulation is given. This largely follows the derivation from Grib et al. (1994, 1972). In Sec. III, we present expressions for the particle momentum distribution function using the exact analytical solution for the mode function in the case of a Sauter-pulsed electric field. Results are discussed in Sec. IV. The article is concluded in Sec. V.
Throughout the paper, we use natural units and set $\hslash=c=m=|e|=1$, and express all variables in terms of the electron mass unit.
@@ II Theory
where $A^{\mu}$ is the four-vector potential, $\Psi(\bm{x},t)$ is the four-component spinor. For Dirac gamma matrices $\gamma^{\mu}$ we chose the Weyl basis Ryder (1996)
@@
and inserting Eq. (5) into Eq. (1), it follows that $\chi(\bm{x},t)$ satisfies the quadratic Dirac equation
@@
$[\partial_{\mu}\partial^{\mu}+e^{2}A^{2}(t)+2iA(t)\partial_{3}-ie\partial_{t}A(t)\gamma^{0}\gamma^{3}+m^{2}]\chi(\bm{x},t)=0.$ (7)
@@
There are two doubly degenerate eigenvectors for $\gamma^{0}\gamma^{3}$: $R_{1}$ and $R_{2}$ with eigenvalue ``$1"$, and $R_{3}$ and $R_{4}$ with eigenvalue ``$-1"$. However, selecting one from each pair suffices Grib et al. (1994). Now, we will seek the solutions of Eq. (9) in the form
@@
Let’s examine the resulting solutions. According to Eq. (14), in a vanishing electric field as $t\rightarrow-\infty$, $\omega(\bm{p},t)$ becomes independent of time, denoted $\omega(\bm{p})=\sqrt{m^{2}+\bm{p}^{2}}$ The scalar function $\psi(\bm{p},t)$ then satisfies the asymptotic equation,
@@
There are two linearly independent solutions to this time-dependent harmonic oscillator equation, which correspond to energy ±$\pm\omega({\bm{p}})$. In what follows, we will label these solutions with superscripts $\lambda=+$ and $\lambda=-$, respectively. The solutions to Eq. (15), are clearly given by plane waves. Therefore,
@@
$\sum_{\lambda=\pm}\sum_{r=\pm}\int\frac{d^{3}{\bm{p}}}{(2\pi)^{3}}\,\Psi_{{\bm{p}}r}^{(\lambda)}(\bm{x},t)[\Psi_{{\bm{p}}r}^{(\lambda)}(x^{\prime},t)]^{\dagger}=\delta({\bm{x}}-{\bm{x}}^{\prime}).$ (20)
@@
$\phi_{\bm{p}}^{(\lambda)}(t)=\frac{e^{-\textrm{i}\lambda\Theta_{\bm{p}}(t_{0},t)}}{\sqrt{2\omega(\bm{p},t)\bigl{(}\omega(\bm{p},t)-\lambda P(p_{\parallel},t)\bigr{)}}},$ (35)
@@
[…]
The lower limit $t_{0}$ of the integral in Eq. (36) is not fully determined as it only fixes an arbitrary phase. Note that the functions $\phi_{\bm{p}}^{(\pm)}(t)$ are chosen such that they coincide with the mode functions $\psi_{\bm{p}}^{(\pm)}(t)$ in the case of a vanishing electric field Smolyansky et al. (1998); Kluger:1998bm.
@@
From the above equations, if we know $\psi_{\bm{p}}(t)$ from a solution of differential Eq. (14) for a specific electric field, we can find out the Bogolyubov transformation coefficients.
@@
In the quasi-particle representation, $f_{r}(\bm{p},t)$ and $\bar{f}_{r}(-\bm{p},t)$ will act as single-particle distribution functions Schmidt et al. (1998). Since the Hamiltonian Eq.(24) does not have any spin-dependent terms, the spin index$(r)$ is dropped, and the distribution function
@@
where the factor ``$2"$ corresponds to the spin degree of freedom. One approach to find $f(\bm{p},t)$ involves solving Eq. (14) for $\psi_{\bm{p}}(t)$ and subsequently using it to determine the Bogoliubov coefficients, which are essential for obtaining the single-particle distribution function.

### v6

@@ Abstract
We investigate the finite-time behavior of pair production from the vacuum by a time-dependent Sauter-pulsed electric field. By examining the temporal behavior of the single-particle distribution function, we observe oscillatory patterns in the longitudinal momentum spectrum at finite times. These oscillations arise due to quantum interference effects resulting from the various dynamical processes/channels leading to the creation of the (quasi-)particle with a given momentum. Furthermore, we derive a simplified approximate analytical expression for the distribution function at finite times, which allows us to explain the origin and behavior of these oscillations. The role of the vacuum polarization function and its counterterm is also discussed in this regard. At finite times, the transverse momentum spectrum peaks at a nonzero value of transverse momentum, indicating the role of multiphoton transitions in quasiparticle creation. We explicitly relate the quasiparticle distribution function to the measurable probability of real pair production when the field is switched off at a finite time, establishing the physical relevance of the observed finite-time structures. This work bridges the quasiparticle description with operationally measurable quantities, offering insights relevant for upcoming high-intensity laser experiments aimed at observing nonperturbative pair production phenomena.
@@ I Introduction
The concept of pair production in an electromagnetic field has its roots in the mid-1920s after the invention of quantum mechanics, with the formulation of the relativistic wave equation for electrons by Paul Dirac in 1928 Dirac (1928). The Dirac Sea model was proposed to explain the enigma of negative energy solutions. F. Sauter’s work in 1931 demonstrated that strong electric fields can lead to pair creation through tunneling with exponential suppression Sauter (1931). This paved the way for quantum field theory, recognizing the vacuum as a polarizable medium influenced by constant fluctuations. In 1936, W. Heisenberg and H. Euler further explored the peculiarities of the Dirac equation, revealing non-linear modifications in Maxwell’s equations due to the interaction of electromagnetic fields with the electron vacuum loop Heisenberg and Euler (1936). J. Schwinger’s groundbreaking work in 1951 precisely calculated the imaginary part of the one-loop effective Lagrangian in the presence of a static electric field Schwinger (1951). As a result of his seminal work, the phenomenon of vacuum pair creation by electric fields has since become widely known as the Schwinger effect, and it is also famously referred to as the Sauter-Schwinger effect in recognition of F. Sauter’s prior work on solving the Dirac equation in the presence of an electric field Sauter (1932).
@@
Schwinger’s pioneering calculation opened up new avenues of research in quantum field theory and has profoundly impacted our understanding of particle physics in the presence of strong fields. This extraordinary property of quantum vacuum producing spontaneous particle-antiparticle pairs has far-reaching implications for understanding the generation of particle-antiparticle pairs in the presence of a strong electric field Greiner et al. (1985); particle creation in the expanding universeParker (1968); black hole evaporation as a result of Hawking radiation Hawking (1975); Parikh and Wilczek (2000); and Unruh radiation, in which particle production is seen by an accelerating observerUnruh (1976); Kim (2016). The study of generating electron-positron pairs through a spatially constant electrical background field was extended to the electric field with various time dependences. In the 1970s, researchers explored the occurrence of pair production from the vacuum in the presence of an oscillating time-dependent electric field Popov (1971); Marinov and Popov (1977). Their investigations revealed different qualitative behaviors for this process under various interaction regimes. The interaction regimes can be distinguished by the value of the dimensionless Keldysh parameter, $\gamma=\frac{m\omega}{|e|E_{0}}$, where $E_{0}$ is the field amplitude, $\omega$ is the field frequency, $e$ is the electron charge, and $m$ is the electron mass Keldysh (1965). When $\gamma>>1$, the process probability exhibits perturbative power-law scaling with field intensity. Instead, for $\gamma<<1$, it exhibits a manifestly non-perturbative exponential dependence on $\frac{1}{E_{0}}$, similar to the case of a constant electric field Nikishov (1969).
@@
Nevertheless, the study of pair production in strong electric fields has attracted sustained interest from theoreticians in recent years due to the extraordinary progress in the development of the ultra-intense lasers technique and the strong-field QED experimental studies that are planned at upcoming high-intensity laser facilities, such as the European X-Ray Free-Electron Laser Andruszkow and others (2000), the Extreme-Light Infrastructure Heinzl and Ilderton (2009); Dunne (2009), the Exawatt Center for Extreme Light Studies Khazanov and et al. (2023). The electric field strength is fast getting closer to the critical value $E_{\text{c}}$.
@@
Various methodologies have been developed to investigate pair production in the presence of strong fields. These include the proper-time method Schwinger (1951); DeWitt (1975), the canonical method Mottola (1985); Tomaras et al. (2001, 2000), Green’s function techniques Gavrilov and Gitman (1996), semiclassical tunneling Parikh and Wilczek (2000), the Schrödinger-functional approach Hallin and Liljenberg (1995), functional techniques Fried and Woodard (2002); Avan et al. (2003); Ilderton (2014), the Wigner formalismZhuang and Heinz (1996); Hebenstreit et al. (2011b, a),mean-field treatments Smolyansky et al. (1998), and worldline instanton methods Affleck et al. (1982); Kim and Page (2002). Studies on pair production in intense laser fields Alkofer et al. (2001); Roberts et al. (2002) have employed analytical and numerical approaches to predict pair-production rates. However, these investigations primarily focus on pair formation at asymptotic times, which alone is insufficient for fully capturing the dynamical aspects of pair production and its time-dependent effects. Several authors have derived expressions for the number of pairs created after a time $T$, where $T$ exceeds the duration of the electric field pulse or corresponds to a moment after the field has vanished Gavrilov and Gitman (1996); Di Piazza (2004); Melike et al. (2012); Ilderton (2022); Adorno et al. (2017). These expressions provide a reliable approximation of the actual pair count.
@@
Understanding pair production dynamics from the initial to final states requires a framework that accounts for transient and non-equilibrium effects. In this context, the distribution function of quasiparticles, governed by the quantum kinetic equation (QKE), offers a suitable description of the time-dependent evolution of pair production Blaschke et al. (2015); Smolyansky et al. (1998). Defining the particle distribution function in the quasiparticle basis is both physically meaningful and intuitive. This approach is grounded in the straightforward physics of how a time-dependent field distorts the time-invariance symmetry of the Hamiltonian, leading to mixing between momentum states of positive and negative energies. This method is particularly relevant for exploring the complex dynamical behavior of pair production in spatially uniform, time-dependent electric fields. Specifically, it has been demonstrated that the temporal evolution of the quasiparticle distribution function in momentum space provides insights into the asymptotic states of the quantum field and offers a comprehensive description of the process during the interaction between the matter field and the strong field background Blaschke et al. (2017). It was shown that the distribution function evolves through three distinct temporal stages: the quasi-electron-positron plasma (QEPP) stage, the transient stage, and finally the residual electron-positron plasma (REPP) stage Banerjee and Singh (2022, 2019). In examining the pair production process, both the QEPP stage, where quasi-particles are formed, and the REPP stage, where real particles emerge, are crucial Blaschke et al. (2013); Shen et al. (2018). While the Schwinger mechanism is primarily known for producing elementary particles in a vacuum under external fields, it can also generate quasiparticle excitations in exotic materials such as graphene Allor et al. (2008); Zubkov (2012); Gavrilov et al. (2012); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015), which are experimentally more accessible, and in superconductors Solinas et al. (2021). Moreover, while ultracold atoms in optical lattices have been investigated as simulators for the Schwinger mechanism Kasper et al. (2016); Halimeh et al. (2025).
@@
In the present work, we consider the production of electron-positron pairs from the vacuum in a time-varying, spatially uniform pulsed electric field given by $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, with amplitude $E_{0}$ and pulse duration $\tau$. Such background field has received extensive attention in the literature Narozhnyi and Nikishov (1970); Dunne and Hall (1998); Balantekin et al. (1991); Gavrilov and Gitman (1996); Nikishov (2004); Levai and Skokov (2010); Klimchitskaya and Mostepanenko (2013), with a focus on the asymptotic behavior of the probability of pair production. To this end, we study the evolution of the single-particle distribution function, $f(\bm{p},t)$, which is rigorously derived by canonical quantization of the Dirac field and subsequent Bogoliubov transformation to a quasi-particle representation Smolyansky et al. (2017).
@@
The distribution function exhibits three distinct dynamical stagesBanerjee and Singh (2022); Otto et al. (2016); Nousch et al. (2017); Smolyansky et al. (2017). By analyzing the temporal evolution of $f(\bm{p},t)$, we find that the emergence and characteristics of these dynamical stages, QEPP, transient, and REPP, are qualitatively and quantitatively influenced by both longitudinal and transverse components of momentum. Next, we analyze the time evolution of the longitudinal momentum spectrum (LMS) of created particles. In the tunneling regime ($\gamma<1$), we observe oscillatory behavior in the LMS at time $t>2\tau$, and this oscillation pattern continuously changes from $t>2\tau$ to $t<6\tau$. This oscillatory behavior at finite times clearly illustrates the quantum interference effects arising from various dynamical processes and channels contributing to the creation of (quasi-)particles with a given momentum. It can also be understood in terms of the vacuum polarization function $u(\bm{p},t)$ and its counterpart, the depolarization function $v(\bm{p},t)$, which encapsulate the underlying particle production dynamics. We emphasize that the oscillations seen in the LMS are not artifacts but rather possess significant physical relevance. In the multi-photon regime, we observe that LMS at finite times near $t=3\tau$ exhibits a multi-modal structure.
@@
Furthermore, we develop an analytical framework valid for finite times $t>\tau$. Specifically, we derive an analytical expression for the single-particle distribution function as a power series in the small parameter (1y)(1-y), where $y=\frac{1}{2}(1+\tanh{(\frac{t}{\tau})})$. The intriguing dynamical features of the momentum distribution function at finite times arise from a function appearing in this expansion. The finite-time behavior of the momentum spectra can be understood in terms of functions that reveal three distinct behaviors, which are discussed in detail in the results section.
@@
Additionally, we explored how the transverse momentum affects LMS and discovered that the oscillatory behavior depends on the value of transverse momentum at that time. Subsequently, we analyzed the temporal progression of the transverse momentum spectrum(TMS) of generated pairs, providing valuable insights into the pair production process. At finite times, the transverse momentum spectra reach their peak at a nonzero value, reflecting the influence of multi-photon transitions in quasiparticle creation.
Crucially, following the approach of Ref. Ilderton (2022), we relate the quasiparticle distribution function $f(\bm{p},t)$ to an experimentally meaningful quantity: the probability $\mathcal{W}(\bm{p},T)$ of producing real electron-positron pairs when the electric field is switched off at a finite time $t=T$. We demonstrate explicitly that $f(\bm{p},T)=\mathcal{W}(\bm{p},T)$, thereby bridging the quasiparticle description with an operationally measurable particle count. This equivalence ensures that the oscillatory structures and momentum spectra analyzed in this work correspond to observable features in pulsed-field experiments.
@@
This article is organized as follows: In Sec. II detailed the theoretical formulation is presented. This largely follows the derivation from Grib et al. (1994, 1972). In Sec. III, we present expressions for the particle momentum distribution function using the exact analytical solution for the mode function in the case of a Sauter-pulsed electric field. Results are discussed in Sec. IV. In Sec. V, we provide a physical interpretation linking the quasiparticle distribution to measurable pair production. The article is concluded in Sec. VI.
Throughout the paper, we use natural units and set $\hslash=c=m=1$, and express all variables in terms of the electron mass units.
@@ II Theory
where $A^{\mu}$ is the four-vector potential, $\Psi(\bm{x},t)$ is a four-component spinor. For Dirac gamma matrices $\gamma^{\mu}$ we chose the Weyl basis Ryder (1996)
@@
and inserting Eq. (5) into Eq. (1), it follows that $\chi(\bm{x},t)$ satisfies the second-order Dirac equation
@@
$\left[\partial_{\mu}\partial^{\mu}+e^{2}A^{2}(t)+2iA(t)\partial_{3}-ie\partial_{t}A(t)\gamma^{0}\gamma^{3}+m^{2}\right]\chi(\bm{x},t)=0.$ (7)
@@
There are two doubly degenerate eigenvectors for $\gamma^{0}\gamma^{3}$: $R_{1}$ and $R_{2}$ with eigenvalue 1``1", and $R_{3}$ and $R_{4}$ with eigenvalue 1``-1". However, selecting one from each pair suffices Grib et al. (1994). Now, we will seek the solutions of Eq. (9) in the form
@@
Let’s examine the resulting solutions. According to Eq. (14), in a vanishing electric field as $t\rightarrow$-$\infty$, $\omega(\bm{p},t)$ becomes independent of time, denoted $\omega(\bm{p})=\sqrt{m^{2}+\bm{p}^{2}}$ . The scalar function $\psi_{\bm{p}}(t)$ then satisfies the asymptotic equation,
@@
There are two linearly independent solutions to this time-dependent harmonic oscillator equation, which correspond to energy $\pm\omega({\bm{p}})$. In what follows, we will label these solutions with superscripts $\lambda=+$ and $\lambda=$-, respectively. The solutions to Eq. (15), are clearly given by plane waves. Therefore,
@@
$\sum_{\lambda}\sum_{r}\int\frac{d^{3}{\bm{p}}}{(2\pi)^{3}}\,\Psi_{{\bm{p}}r}^{(\lambda)}(\bm{x},t)[\Psi_{{\bm{p}}r}^{(\lambda)}(\bm{x}^{\prime},t)]^{\dagger}=\delta({\bm{x}}-{\bm{x}}^{\prime}).$ (20)
@@
$\phi_{\bm{p}}^{(\lambda)}(t)=\frac{e^{-\textrm{i}\lambda\Theta_{\bm{p}}(t_{0},t)}}{\sqrt{2\omega(\bm{p},t)\bigl[\omega(\bm{p},t)-\lambda P(p_{\parallel},t)\bigr]}},$ (35)
@@
$\Theta_{\bm{p}}(t_{0},t)$ $=\int^{t}_{t_{0}}\omega(\bm{p},t^{\prime})dt^{\prime}.$ (36)
The lower limit $t_{0}$ of the integral in Eq. (36) is not fully determined as it only fixes an arbitrary phase. Note that the functions $\phi_{\bm{p}}^{(\pm)}(t)$ are chosen such that they coincide with the mode functions $\psi_{\bm{p}}^{(\pm)}(t)$ in the case of a vanishing electric field Smolyansky et al. (1998); Kluger et al. (1998).
@@
From the above equations, if we know $\psi_{\bm{p}}(t)$ from a solution of differential Eq. (14) for a specific electric field, we can find out the Bogoliubov transformation coefficients.
@@
In the quasi-particle representation, $f_{r}(\bm{p},t)$ and $\bar{f}_{r}(-\bm{p},t)$ will act as single-particle distribution functions Schmidt et al. (1998). Since the Hamiltonian Eq.(24) does not have any spin-dependent terms, the spin index(r)(r) is dropped, and the distribution function
@@
where the factor 2``2" corresponds to the spin degree of freedom. One approach to find $f(\bm{p},t)$ involves solving Eq. (14) for $\psi_{\bm{p}}(t)$ and subsequently using it to determine the Bogoliubov coefficients, which are essential for obtaining the single-particle distribution function.